# Upper Approximation Bounds for Neural Oscillators

### v1

@@ A.4 Proof of Lemma 4
$\begin{array}[]{l}\sum\limits_{n=1}^{+\infty}{\left|{{\omega_{n}}}\right|\left|{{\alpha_{n}}}\right|}=\frac{1}{T}\sum\limits_{n=1}^{+\infty}{\left|{\int_{-2T}^{2T}{{\omega_{n}}{e^{-{\rm{i}}{\omega_{n}}\tau}}{\rho_{v}}(\tau){\rm{d}}\tau}}\right|}\mathop{\leq}\limits^{(b)}\frac{1}{T}\sum\limits_{n=1}^{+\infty}{\frac{1}{{\omega_{n}^{2}}}\int_{-v}^{0}{\left|{{{\rho}^{\prime\prime\prime}_{v}}(\tau)}\right|{\rm{d}}\tau}}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\leq\frac{1}{T}\sum\limits_{n=1}^{+\infty}{\frac{4{T^{2}}v}{{\pi^{2}}{n^{2}}}}\mathop{\max}\limits_{t\in[-v,0]}\left|{{{\rho}^{\prime\prime\prime}_{v}}(t)}\right|\leq\frac{{4T{c_{\rho,3}}}}{{{\pi^{2}}{v^{3}}}}\sum\limits_{n=1}^{+\infty}{\frac{1}{{{n^{2}}}}}\mathop{=}\limits^{(c)}\frac{{{\pi^{2}}M{c_{\rho,3}}}}{{24T^{2}{c_{\rho,2}}}},\end{array}$ (70)
Step (b) is derived using the same method as that in Eq. (45), Step (c) holds from Eq. (14) and $\sum_{n=1}^{+\infty}{n^{-2}}=\pi^{2}/6$, and $c_{\rho,3}=\max_{\tau\in[-1,0]}\left|{\rho}_{1}^{\prime\prime\prime}(\tau)\right|$ is an independent constant with ${\rho}_{1}(\tau)$ in Eq. (41).
@@
$\begin{array}[]{l}\left|{{B_{n}}(\Delta t)}\right|\\ \;\;\;\leq\left|{\cos\left[{{\omega_{n}}(t+\Delta t)}\right]\sin\left[{{\omega_{n}}(t+\Delta t-\tau)-{\theta_{n}}}\right]-\cos\left[{{\omega_{n}}(t+\Delta t)}\right]\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]}\right|\\ \;\;\;\;\;\;\;+\left|{\cos\left[{{\omega_{n}}(t+\Delta t)}\right]\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]-\cos({\omega_{n}}t)\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]}\right|\leq 2\left|{{\omega_{n}}}\right|\left|{\Delta t}\right|.\end{array}$ (71)
@@
$\left|{{{{\bf{\hat{u}}}}_{t+\Delta t}}(\tau)-{{{\bf{\hat{u}}}}_{t}}(\tau)}\right|\leq p\left|{\Delta t}\right|{B_{K}}MT^{-1}\left({\frac{{{\pi^{2}}{c_{\rho,3}}}}{{24{c_{\rho,2}}}}+3}\right).$ (73)
@@
$\left|{{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t+\Delta t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t)}\right|\leq\left|{\Delta t}\right|{D_{{\Phi_{j}}}}+{\phi_{{\Phi_{j}}}}\left(p{\left|{\Delta t}\right|{B_{K}}MT^{-1}c_{L}}\right),$ (74)
where $c_{L}=24^{-1}{\pi^{2}}{c_{\rho,3}}{c_{\rho,2}^{-1}}+3$ is an independent constant. The first-order derivative of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ with respect to t is bounded by
$\begin{array}[]{l}\left|{\frac{{{\partial}{\psi_{j}}({\mathbf{x}},{{\bf{u}}_{0}},t)}}{{{\partial}t}}}\right|\leq{D_{{\Phi_{j}}}}+\mathop{\lim}\limits_{\left|{\Delta t}\right|\to 0}{\left|{\Delta t}\right|^{-1}}{\phi_{{\Phi_{j}}}}\left({p\left|{\Delta t}\right|{B_{K}}{M}T^{-1}c_{L}}\right)={D_{{\Phi_{j}}}}+p{B_{K}}MT^{-1}c_{L}{{\phi^{\prime}}_{{\Phi_{j}}}}(0).\end{array}$ (75)
@@
$\begin{array}[]{l}\left|{{\psi_{j}}({\bf{x}}+\Delta{\bf{x}},{{\bf{u}}_{0}}+\Delta{{\bf{u}}_{0}},t+\Delta t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t)}\right|\\ \;\;\;\leq\left|{{\psi_{j}}({\bf{x}}+\Delta{\bf{x}},{{\bf{u}}_{0}}+\Delta{{\bf{u}}_{0}},t+\Delta t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}}+\Delta{{\bf{u}}_{0}},t+\Delta t)}\right|\\ \;\;\;\;\;\;\;+\left|{{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}}+\Delta{{\bf{u}}_{0}},t+\Delta t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t+\Delta t)}\right|+\left|{{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t+\Delta t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t)}\right|\\ \;\;\;\leq{L_{{\psi_{j}}}\left(M\right)}\left({\left|{\Delta{\bf{x}}}\right|+\left|{\Delta{{\bf{u}}_{0}}}\right|+\left|{\Delta t}\right|}\right)\mathop{\leq}\limits^{(a)}\sqrt{p(M+1)+1}{L_{{\psi_{j}}}\left(M\right)}\sqrt{\left|{\Delta{\bf{x}}}\right|_{2}^{2}+\left|{\Delta{{\bf{u}}_{0}}}\right|_{2}^{2}+\left|{\Delta t}\right|_{2}^{2}},\end{array}$ (76)
@@ A.5 Proof of Theorem 1
Proof: From Remark 2 and Lemma 2, for an arbitrary positive integer $M_{\it{\Gamma}}$ satisfying Eq. (15) along with v satisfying Eq. (14), there exists a one-hidden-layer MLP ${\it{\Gamma}}(\cdot)$ specified in Lemma 2, for every $\mathbf{u}(t)\in K$, its corresponding solution $\mathbf{x}(t)=[\mathbf{x}_{1}^{\top}(t),\mathbf{x}_{2}^{\top}(t),...,\mathbf{x}_{M_{\it{\Gamma}}}^{\top}(t)]^{\top}$ calculated from the second-order ODE in Eq. (1) with the initial conditions $\mathbf{x}(0)={\mathbf{x}}^{\prime}(0)=\mathbf{0}$ equal the sine transform coefficients of $\mathbf{u}(t)$, as shown in Eq. (55). From Lemma 3 and Remark 3, there exists a continuous mapping $\Psi=[\psi_{1},\psi_{2},...,\psi_{q}]^{\top}:\mathbb{R}^{p\left(M_{\it{\Gamma}}+1\right)}\times\left[0,T\right]\to\mathbb{R}^{q}$ for all $\mathbf{u}(t)\in K$, the difference between $\Phi[\mathbf{u}(\tau)](t)$ and $\Psi[\mathbf{x}(t),\mathbf{u}(0),t]$ can be bounded by
$\left|{\Phi\left[{{\bf{u}}(\tau)}\right](t)-\Psi\left[{{\bf{x}}(t),{\bf{u}}(0),t}\right]}\right|\leq{\phi_{\Phi}}\left(\frac{{54^{1/3}p{L_{K}}Tc_{\rho,2}^{{1\mathord{\left/{\vphantom{13}}\right.\kern-1.2pt}3}}}}{{{\pi^{{2\mathord{\left/{\vphantom{23}}\right.\kern-1.2pt}3}}}{M_{{\it{\Gamma}}}^{{1\mathord{\left/{\vphantom{13}}\right.\kern-1.2pt}3}}}}}\right).$ (77)
From Lemma 4, the continuous function $\Psi\left({\bf{x}},{\bf{u}}_{0},t\right)$ in Eq. (77) is compactly supported within $\Omega_{M_{\it{\Gamma}}}\left(TB_{K},B_{K},T\right)=\left[-TB_{K},TB_{K}\right]^{pM_{\it{\Gamma}}}\times\left[-B_{K},B_{K}\right]^{p}\times\left[0,T\right]$ for every integer $M_{\it{\Gamma}}$ satisfying Eq. (15) and all $\mathbf{u}(t)\in K$, where $\mathbf{x}$ and $\mathbf{u}_{0}$ represent the values of $\mathbf{x}(t)$ and $\mathbf{u}(0)$, respectively. The Lipschtiz constant $L_{{\psi_{j}}}\left(M_{\it{\Gamma}}\right)$ of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ of $\Psi\left({\bf{x}},{\bf{u}}_{0},t\right)$ with respect to its each argument is in Eq. (17) and a linear modulus of continuity $\phi_{\psi_{j},M_{{\it{\Gamma}}}}(x)$ of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ under $L^{2}$-norm is in Eq. (18).
Theorem 1 by Hanin (2019) demonstrates that given an arbitrary continuous function $f\left(\mathbf{z}\right)$, $\mathbf{z}=\left[z_{1},z_{2},...,z_{d}\right]^{\top}$, mapping from $0,1]^{d}$ to $[0,1]$, with a modulus of continuity $w_{f}(z)$ of $f\left(\mathbf{z}\right)$, for an arbitrary error $\varepsilon>0$, there exists an MLP $f_{N}\left(\mathbf{z}\right)$ with $\sigma_{\mathrm{ReLU}}(\cdot)$, input dimension $d$, hidden layer width $d+3$, output dimension 1, and depth $2d!\left[\sqrt{d}/w_{f}^{-1}(\varepsilon)\right]^{d}$, where $w_{f}^{-1}(\varepsilon)$ is the inverse function of $w_{f}\left(z\right)$, such that $\sup_{\mathbf{z}\in\left[0,1\right]^{d}}\left|f\left(\mathbf{z}\right)-f_{N}\left(\mathbf{z}\right)\right|\leq\varepsilon$. Theorem 1 by Hanin and Sellke (2017) and Proposition 53 by Kratsios and Papon (2022) provide similar results.
@@ B.1 Proof of Lemma 5
$\begin{array}[]{l}{{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})\\ \;\;\;\;\;=\int_{0}^{t}{\frac{{{\rm{d}}{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{{\rm{d}}\tau}}{\rm{d}}\tau}=\int_{0}^{t}{\frac{{\rm{d}}}{{{\rm{d}}\tau}}\left\{{\int_{\tau}^{t}{{{{\bf{z}}}_{1}^{\prime}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]{\rm{d}}\lambda}+{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right\}{\rm{d}}\tau}\\ \;\;\;\;\;=\int_{0}^{t}{\left\{{\int_{\tau}^{t}{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial\tau}}{\rm{d}}\lambda}+\frac{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{{\partial\tau}}-{{\left.{\frac{{\partial{{\bf{z}}_{1}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial\lambda}}}\right|}_{\lambda=\tau}}}\right\}{\rm{d}}\tau}\\ \;\;\;\;\;=\int_{0}^{t}{\left({\left\{{\int_{\tau}^{t}{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{\rm{d}}\lambda}+{\bf{I}}}\right\}\frac{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{{\partial\tau}}-{{\left.{\frac{{\partial{{\bf{z}}_{1}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial\lambda}}}\right|}_{\lambda=\tau}}}\right){\rm{d}}\tau}\\ \;\;\;\;\;=\int_{0}^{t}{{\left\{{\int_{\tau}^{t}{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{\rm{d}}\lambda}+{\bf{I}}}\right\}\par{\begin{Bmatrix}{{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}})}\\ {{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\end{Bmatrix}}}{\rm{d}}\tau}-\int_{0}^{t}{{{{\left.{\frac{{\partial{{\bf{z}}_{1}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial\lambda}}}\right|}_{\lambda=\tau}}}{\rm{d}}\tau},\end{array}$ (80)
@@
[…]

### v2

@@ A.4 Proof of Lemma 4
$\sum\limits_{n=1}^{+\infty}{\left|{{\omega_{n}}}\right|\left|{{\alpha_{n}}}\right|}\mathop{=}\limits^{(b)}\frac{1}{T}\sum\limits_{n=1}^{+\infty}\left|{\omega_{n}}{\int_{-2T}^{2T}{{e^{-{\rm{i}}{\omega_{n}}\tau}}{\rho_{v}}(\tau){\rm{d}}\tau}}\right|\leq\frac{1}{T}\sum\limits_{n=1}^{+\infty}{\frac{1}{{\omega_{n}^{2}}}\int_{-v}^{0}{\left|{{{\rho}^{\prime\prime\prime}_{v}}(\tau)}\right|{\rm{d}}\tau}}$ (75)
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\leq\frac{1}{T}\sum\limits_{n=1}^{+\infty}{\frac{4{T^{2}}v}{{\pi^{2}}{n^{2}}}}\mathop{\max}\limits_{t\in[-v,0]}\left|{{{\rho}^{\prime\prime\prime}_{v}}(t)}\right|\leq\frac{{4T{c_{\rho,3}}}}{{{\pi^{2}}{v^{3}}}}\sum\limits_{n=1}^{+\infty}{\frac{1}{{{n^{2}}}}}\mathop{=}\limits^{(c)}\frac{{2c_{\rho,3}M^{3}}}{3T^{2}},$
Step (b) is from Eq. (56), Step (c) holds from Eq. (15) and $\sum_{n=1}^{+\infty}{n^{-2}}=\pi^{2}/6$, and $c_{\rho,3}=\max_{\tau\in[-1,0]}\left|{\rho}_{1}^{\prime\prime\prime}(\tau)\right|$ is an independent constant with ${\rho}_{1}(\tau)$ in Eq. (46).
@@
$\left|{{B_{n}}(\Delta t)}\right|\leq\left|{\cos\left[{{\omega_{n}}(t+\Delta t)}\right]\sin\left[{{\omega_{n}}(t+\Delta t-\tau)-{\theta_{n}}}\right]-\cos\left[{{\omega_{n}}(t+\Delta t)}\right]\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]}\right|$ (76)
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\left|{\cos\left[{{\omega_{n}}(t+\Delta t)}\right]\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]-\cos({\omega_{n}}t)\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]}\right|\leq 2\left|{{\omega_{n}}}\right|\left|{\Delta t}\right|.$
@@
$\left|{{{{\bf{\hat{u}}}}_{t+\Delta t}}(\tau)-{{{\bf{\hat{u}}}}_{t}}(\tau)}\right|\leq\left|{\Delta t}\right|p{B_{K}}M^{3}T^{-1}\left({c_{\rho,3}+3}\right).$ (78)
@@
$\left|{{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t+\Delta t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t)}\right|\leq\left|{\Delta t}\right|{D_{{\Phi_{j}}}}+{\phi_{{\Phi_{j}}}}\left({\left|{\Delta t}\right|p{B_{K}}M^{3}T^{-1}c_{L}}\right),$ (79)
where $c_{L}=c_{\rho,3}+3$ is an independent constant. The first-order derivative of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ with respect to t is bounded by
$\left|{\frac{{{\partial}{\psi_{j}}({\mathbf{x}},{{\bf{u}}_{0}},t)}}{{{\partial}t}}}\right|\leq{D_{{\Phi_{j}}}}+\mathop{\lim}\limits_{\left|{\Delta t}\right|\to 0}{\left|{\Delta t}\right|^{-1}}{\phi_{{\Phi_{j}}}}\left({\left|{\Delta t}\right|p{B_{K}}{M^{3}}T^{-1}c_{L}}\right)={D_{{\Phi_{j}}}}+p{B_{K}}M^{3}T^{-1}c_{L}{{\phi^{\prime}}_{{\Phi_{j}}}}(0).$ (80)
@@
$\left|{{\psi_{j}}({\bf{x}}+\Delta{\bf{x}},{{\bf{u}}_{0}}+\Delta{{\bf{u}}_{0}},t+\Delta t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t)}\right|$ (81)
$\;\;\;\;\;\;\leq\left|{{\psi_{j}}({\bf{x}}+\Delta{\bf{x}},{{\bf{u}}_{0}}+\Delta{{\bf{u}}_{0}},t+\Delta t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}}+\Delta{{\bf{u}}_{0}},t+\Delta t)}\right|$
$\;\;\;\;\;\;\;\;\;\;+\left|{{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}}+\Delta{{\bf{u}}_{0}},t+\Delta t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t+\Delta t)}\right|+\left|{{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t+\Delta t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t)}\right|$
$\;\;\;\;\;\;\leq{L_{{\psi_{j}}}\left(M\right)}\left({\left|{\Delta{\bf{x}}}\right|+\left|{\Delta{{\bf{u}}_{0}}}\right|+\left|{\Delta t}\right|}\right)\mathop{\leq}\limits^{(a)}\sqrt{p(M+1)+1}{L_{{\psi_{j}}}\left(M\right)}\sqrt{\left|{\Delta{\bf{x}}}\right|_{2}^{2}+\left|{\Delta{{\bf{u}}_{0}}}\right|_{2}^{2}+\left|{\Delta t}\right|_{2}^{2}},$
@@ A.5 Proof of Theorem 1
Proof: From Remark 2 and Lemma 2, for an arbitrary integer $M_{\it{\Gamma}}$ satisfying Eq. (15) along with v satisfying Eq. (14), there exists a one-hidden-layer MLP ${\it{\Gamma}}(\cdot)$ specified in Lemma 2, for every $\mathbf{u}(t)\in K$, its corresponding solution $\mathbf{x}(t)=[\mathbf{x}_{1}^{\top}(t),\mathbf{x}_{2}^{\top}(t),...,\mathbf{x}_{M_{\it{\Gamma}}}^{\top}(t)]^{\top}$ calculated from the second-order ODE in Eq. (1) with the initial conditions $\mathbf{x}(0)={\mathbf{x}}^{\prime}(0)=\mathbf{0}$ equals the sine transform coefficients of $\mathbf{u}(t)$, as shown in Eq. (60). From Lemma 3 and Remark 3, there exists a continuous mapping $\Psi=[\psi_{1},\psi_{2},...,\psi_{q}]^{\top}:\mathbb{R}^{p\left(M_{\it{\Gamma}}+1\right)}\times\left[0,T\right]\to\mathbb{R}^{q}$ for all $\mathbf{u}(t)\in K$, the difference between $\Phi[\mathbf{u}(\tau)](t)$ and $\Psi[\mathbf{x}(t),\mathbf{u}(0),t]$ can be bounded by
$\left|{\Phi\left[{{\bf{u}}(\tau)}\right](t)-\Psi\left[{{\bf{x}}(t),{\bf{u}}(0),t}\right]}\right|\leq{\phi_{\Phi}}\left[\frac{32c_{K}TpL_{K}\left(\ln{M}\right)^{2}}{M}\right].$ (82)
From Lemma 4, the continuous function $\Psi\left({\bf{x}},{\bf{u}}_{0},t\right)$ in Eq. (82) is compactly supported within $\Omega_{M_{\it{\Gamma}}}\left(TB_{K},B_{K},T\right)=\left[-TB_{K},TB_{K}\right]^{pM_{\it{\Gamma}}}\times\left[-B_{K},B_{K}\right]^{p}\times\left[0,T\right]$ for every integer $M_{\it{\Gamma}}$ satisfying Eq. (15) and all $\mathbf{u}(t)\in K$, where $\mathbf{x}$ and $\mathbf{u}_{0}$ represent the values of $\mathbf{x}(t)$ and $\mathbf{u}(0)$, respectively. The Lipschitz constant $L_{{\psi_{j}}}\left(M_{\it{\Gamma}}\right)$ of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ of $\Psi\left({\bf{x}},{\bf{u}}_{0},t\right)$ with respect to its each argument is in Eq. (17) and a linear modulus of continuity $\phi_{\psi_{j},M_{{\it{\Gamma}}}}(x)$ of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ under the $L^{2}$-norm is in Eq. (18).
Theorem 1 by Hanin (2019) demonstrates that given an arbitrary continuous function $f\left(\mathbf{z}\right)$, $\mathbf{z}=\left[z_{1},z_{2},...,z_{d}\right]^{\top}$, mapping from $0,1]^{d}$ to $[0,1]$, with a modulus of continuity $w_{f}(z)$ of $f\left(\mathbf{z}\right)$, for an arbitrary error $\varepsilon>0$, there exists an MLP $f_{N}\left(\mathbf{z}\right)$ with $\sigma_{\mathrm{ReLU}}(\cdot)$, input dimension $d$, hidden-layer width $d+3$, output dimension 1, and depth $2d!\left[\sqrt{d}/w_{f}^{-1}(\varepsilon)\right]^{d}$, where $w_{f}^{-1}(\varepsilon)$ is the inverse function of $w_{f}\left(z\right)$, such that $\sup_{\mathbf{z}\in\left[0,1\right]^{d}}\left|f\left(\mathbf{z}\right)-f_{N}\left(\mathbf{z}\right)\right|\leq\varepsilon$. Theorem 1 by Hanin and Sellke (2017) and Proposition 53 by Kratsios and Papon (2022) provide similar results.
@@ B.1 Proof of Lemma 5
${{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})$ (85)
$\;\;\;\;\;=\int_{0}^{t}{\frac{{{\rm{d}}{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{{\rm{d}}\tau}}{\rm{d}}\tau}=\int_{0}^{t}{\frac{{\rm{d}}}{{{\rm{d}}\tau}}\left\{{\int_{\tau}^{t}{{{{\bf{z}}}_{1}^{\prime}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]{\rm{d}}\lambda}+{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right\}{\rm{d}}\tau}$
$\;\;\;\;\;=\int_{0}^{t}{\left\{{\int_{\tau}^{t}{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial\tau}}{\rm{d}}\lambda}+\frac{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{{\partial\tau}}-{{\left.{\frac{{\partial{{\bf{z}}_{1}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial\lambda}}}\right|}_{\lambda=\tau}}}\right\}{\rm{d}}\tau}$
$\;\;\;\;\;=\int_{0}^{t}{\left({\left\{{\int_{\tau}^{t}{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{\rm{d}}\lambda}+{\bf{I}}}\right\}\frac{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{{\partial\tau}}-{{\left.{\frac{{\partial{{\bf{z}}_{1}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial\lambda}}}\right|}_{\lambda=\tau}}}\right){\rm{d}}\tau}$
$\;\;\;\;\;=\int_{0}^{t}{{\left\{{\int_{\tau}^{t}{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{\rm{d}}\lambda}+{\bf{I}}}\right\}{\begin{Bmatrix}{{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}})}\\ {{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\end{Bmatrix}}}{\rm{d}}\tau}-\int_{0}^{t}{{{{\left.{\frac{{\partial{{\bf{z}}_{1}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial\lambda}}}\right|}_{\lambda=\tau}}}{\rm{d}}\tau},$
@@
$0={\left.{\frac{{{\rm{d}}{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{1}}(\hat{\tau},{{\bf{z}}_{\tau}})}\right]}}{{{\rm{d}}\hat{\tau}}}}\right|_{\hat{\tau}=\tau}}={\left.{\frac{{\rm{d}}}{{{\rm{d}}\hat{\tau}}}\left\{{\int_{\hat{\tau}}^{t}{{{{\bf{z}}}_{1}^{\prime}}\left[{\lambda,{{\bf{z}}_{1}}(\hat{\tau},{{\bf{z}}_{\tau}})}\right]{\rm{d}}\lambda}+{{\bf{z}}_{1}}(\hat{\tau},{{\bf{z}}_{\tau}})}\right\}}\right|_{\hat{\tau}=\tau}}$ (86)
$\;\;\;={\left.{\int_{\hat{\tau}}^{t}{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{\lambda,{{\bf{z}}_{1}}(\hat{\tau},{{\bf{z}}_{0}})}\right]}}{{\partial\hat{\tau}}}{\rm{d}}\lambda}}\right|_{\hat{\tau}=\tau}}+{\left.{\frac{{\partial{{\bf{z}}_{1}}(\hat{\tau},{{\bf{z}}_{\tau}})}}{{\partial\hat{\tau}}}}\right|_{\hat{\tau}=\tau}}-{\left.{\frac{{\partial{{\bf{z}}_{1}}\left[{\lambda,{{\bf{z}}_{1}}(\hat{\tau},{{\bf{z}}_{\tau}})}\right]}}{{\partial\lambda}}}\right|_{\lambda=\hat{\tau},\hat{\tau}=\tau}}$
$\;\;\;={\left.{\left\{{\int_{\hat{\tau}}^{t}{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{\lambda,{{\bf{z}}_{1}}(\hat{\tau},{{\bf{z}}_{\tau}})}\right]}}{{\partial{{\bf{z}}_{1}}(\hat{\tau},{{\bf{z}}_{\tau}})}}{\rm{d}}\lambda}+{\bf{I}}}\right\}\frac{{\partial{{\bf{z}}_{1}}(\hat{\tau},{{\bf{z}}_{\tau}})}}{{\partial\hat{\tau}}}}\right|_{\hat{\tau}=\tau}}-{\left.{\frac{{\partial{{\bf{z}}_{1}}\left[{\lambda,{{\bf{z}}_{1}}(\hat{\tau},{{\bf{z}}_{\tau}})}\right]}}{{\partial\lambda}}}\right|_{\lambda=\hat{\tau},\hat{\tau}=\tau}}$
$\;\;\;={\left.{\left\{{\int_{\hat{\tau}}^{t}{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{\lambda,{{\bf{z}}_{1}}(\hat{\tau},{{\bf{z}}_{\tau}})}\right]}}{{\partial{{\bf{z}}_{1}}(\hat{\tau},{{\bf{z}}_{\tau}})}}{\rm{d}}\lambda}+{\bf{I}}}\right\}{\begin{Bmatrix}{{{\bf{z}}_{1,2}}(\hat{\tau},{{\bf{z}}_{\tau}})}\\ {{g_{1}}\left[{{{\bf{z}}_{1,1}}(\hat{\tau},{{\bf{z}}_{\tau}}),{{\bf{z}}_{1,2}}(\hat{\tau},{{\bf{z}}_{\tau}}),{\bf{u}}(\hat{\tau})}\right]}\end{Bmatrix}}}\right|_{\hat{\tau}=\tau}}-{\left.{\frac{{\partial{{\bf{z}}_{1}}\left[{\lambda,{{\bf{z}}_{1}}(\hat{\tau},{{\bf{z}}_{\tau}})}\right]}}{{\partial\lambda}}}\right|_{\lambda=\hat{\tau},\hat{\tau}=\tau}}.$